OpenStudy (anonymous):

Completely factor the expression 245x^3 – 420x^2y + 180xy^2

OpenStudy (anonymous):

245x^3 – 420x^2y + 180xy^2 =5x(49x^2 - 84xy + 36y^2) =5x[(7x)^2 -2(7x)(6y) + (6y)^2] we know the formula, (a-b)^2= a^2-2ab+b^2 comparing (7x)^2 -2(7x)(6y) + (6y)^2 with above formula , we get (7x - 6y)2 putting this in our equation, we get 5x(7x-6y)^2